A person plans to invest three times as much in a Dreyfus Premier Worldwide Growth R account at 4.5% interest as in an Oppenheimer Global Opportunities Y account at 14.9% annual interest. How much should the person invest in each account to earn a total of $426 in one year?
Answer by richwmiller(17219)

We know that the interest for the two accounts is $426
0.045*x+0.149*y=426
We know that the account at 4.5% has 3 times the amount at 14.9%
We substitute for x
0.045*(3*y)+0.149*y=426
We multiply out
(0.135+0.149)*y=426
We combine like terms.
0.284*y=426
Isolate y
y= 1500.00
y=$ 1500.00 at 14.9%
Calculate x
x=3*y
x=$ 4500.00 at 4.5%
Your total for x is $ 4500.00
Your total for y is $ 1500.00
Your total for x and y is $ 6000.00
Total invested $ 4500.00 + $ 1500.00=$ 6000.00
We check
0.045* 4500.00+0.149* 1500.00=426
202.50+ 223.50=426
426.0=426
Since this statement is TRUE and neither x nor y is negative then all is well.